OpenStudy (anonymous):

A pilot flew a 400-mile flight in 2.5 hours flying into the wind. Flying the same rate and with the same wind speed, the return trip took only 2 hours, with a tailwind. What was the speed of the wind?

OpenStudy (anonymous):

help

OpenStudy (caution0):

let the speed of plane=x m/hr speed of wind =y m/hr then (x-y)*(2.5)=400 x-y=400*(2/5)=160 ...(1) (x+y)*2=400 x+y=400/2=200 ...(2) subtract (1) from (2) and find y

OpenStudy (welshfella):

distance forward = distance back = 400 miles speed = distance / time let speed of the plane with no wind = v and speed of the wind = w:- agst the wind v - w = 400 / 2.5 with the wind v + w = 400 / 2 solve this system of equations to find w.

OpenStudy (caution0):

lml we basically have the same thing

OpenStudy (welshfella):

yes